1. Two charges Q1=-5 NC and Q2=-7 nc are separated by 8 cm. Calculate tHe electric potential at point.

2. A research van de Graaf generator has a 2.50 mm diameter sphere with a charge of 6.00 UC on it.

A. What is the potential near its surface?

B. At what distance from its center is the potential 1.50 MV?



1
Expert's answer
2022-03-10T18:01:34-0500

1.

"\varphi=\frac{k(q_1-q_2)}r=225~V,"

2a.

"\varphi=\frac{kq}r=21.6~MV,"

2b.

"R=\frac{kq}{\varphi}=36~mm."